How do you find i^30?

Answer by Alan3354(21588)

How do you find i^30?
-----------
i^2 = -1
i^4 = 1
i^8 = 1
i^12 = 1, etc.
------------
Divide the exponent by 4, and use the remainder
30/4 = 7 rem 2
i^30 = i^2 = -1